OpenStudy (anonymous):

how would I go about solving 3.8/40

OpenStudy (anonymous):

do you have a calculator?

OpenStudy (anonymous):

3.8=38/10 this means (38/10)/40 =38/(40*10) =38/400 =0.095

OpenStudy (anonymous):

great.

OpenStudy (anonymous):

would you mind helping me with a another problem